A projectile is fired in such a way that its horizontal range is equal to three times its maximum height. What is the angle of projection?
1
Expert's answer
2020-01-03T09:21:01-0500
"R=3H"

"\frac{v^2\sin{2\theta}}{g}=2\frac{v^2\sin{\theta}\cos{\theta}}{g}=3
\frac{v^2\sin^2{\theta}}{2g}"

"\tan \theta=\frac{4}{3}"

"\theta=53\degree"